OpenStudy (anonymous):

Partial Fraction Decomposition: Can someone please help me with the Partial Frac Decomp of the following:

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[1\div(s-1)(s+1)^{2} \] also with \[1\div((s)(s+1)^{2})\]

OpenStudy (anonymous):

In that first one, it's only 1 in the numerator,, forgot to put parenthesis. I'm trying to solve differential equations using LaPlace transform, but i think I'm messing up on the PFD,, everything is coming out correctly except my coefficients in my solution.

OpenStudy (anonymous):

The partial fractions I get are: \[\frac{1}{4}[\frac{1}{s-1}-\frac{1}{s+1}-\frac{2}{(s+1)^{2}}]\]and\[\frac{1}{s}-\frac{1}{s+1}-\frac{1}{(s+1)^{2}}\]

OpenStudy (anonymous):

Thank you